# Towards Effective Prompt Stealing Attack against Text-to-Image Diffusion Models

### v1

@@ I Introduction
Responsible Disclosure: We obtained prior consent from both the prompt engineers and the platforms for the scientific use of the displayed prompt showcases and the collected data. They have expressed great interest in and support for our work. Upon acceptance, we will share the research findings with them, along with mitigation methods for such attacks. Additionally, we will open-source the code and the collected dataset for community use.
@@ II-B Prompt-as-a-Service
Despite the remarkable generation ability of T2I models, performing prompt engineering requires a significant amount of effort and cost. For instance, PromptPerfect charges up to \$100 per month for its prompt optimization service [29]. Under these circumstances, Prompt-as-a-Service (PaaS) has emerged as a popular application. Professional prompt engineers optimize prompts with their expertise and sell them on the prompt markets, e.g., PromptBase [26], AIFrog [3], PromptSea [31], PromptHero [28]. These on-sell prompts are sightlessly exhibited with their exquisite showcases, which visually indicate the effects of prompts. After purchasing a prompt, the user gains access to its corresponding textual description, which can be used to generate customized images or increase the number of showcased outputs.
@@ III-A Threat Model
Attacker’s Capability. We make two assumptions about the attacker’s capabilities. (1) The attacker can access the generated showcases of the prompt (denoted as $s$). This is typical in commercial platforms, where prompt sellers display showcase images to attract potential buyers. However, the specific prompt used to generate the showcase remains hidden from all users until they purchase it. A detailed example can be found in Portrait Neon Lights [30]. (2) As some commercial victim models run in a pay-as-you-go mode, which may be expensive, we assume the attacker can use a substitute local text-to-image (T2I) model, referred to as the proxy model. Since the proxy model can be any open-source model, the query budget is 0.
Adversary’s Motivation. The attacker aims to compromise the confidentiality of valuable prompts sold on the markets. A carefully designed prompt consists of subjects and modifiers (as shown in Figure 1). Specifically, subjects are the main body of a prompt that outlines an image’s most important elements, while modifiers detail the style of the subjects and the whole picture. Formally, given a showcase image $s$, an adversary aims to recover the prompt $p$ that is used to create the showcase, including both its subjects and modifiers.
From a utility perspective, stolen prompts can be reused for customization, e.g., replacing subjects or increasing the number of showcases. From a commercial standpoint, as discussed in previous work [43], the motivation behind prompt stealing can be twofold. First, it may infringe on the intellectual property (IP) of high-quality prompts, which is the result of significant effort by prompt engineers. Second, attackers have strong economic incentives: they can avoid payment by using these premium prompts for free and may also redistribute or resell the prompts on other markets for additional profit. This poses a significant threat to the commercial interests of both prompt developers and the platforms hosting them, which is explicitly prohibited by the industrial platform PromptBase [27]. From a privacy protection standpoint, this type of attack poses a significant threat to the confidentiality of prompts, particularly when they contain sensitive or highly classified information.
@@ III-B Attack Requirements
Attack Comparisons. We follow the existing line of work [15, 25, 43], and focus on reversing readable prompts, which could satisfy these two objectives. We observe that certain hard prompt recovery methods could serve as potential solutions. For instance, PH2P [19] leverages gradient signals to optimize a soft prompt, which is then projected back into a hard prompt. However, since semantic coherence is not taken into account during optimization, the recovered prompts often consist of unreadable or nonsensical words. As a result, they fail to meet the aforementioned objectives, limiting the practical utility of the extracted prompts. We list the merits of our Prometheus over the existing works in Table I. Besides static modifiers, we also include dynamic modifiers and feedback from the proxy T2I model, to steal the prompt with higher fidelity.
@@ IV-A Overview
1. Modifier Extraction (Section IV-B). First, we construct a comprehensive and semantic modifier pool to ensure functionality and reusability. We satisfy these requirements by considering two kinds of modifiers: 1) Static modifiers, which are extracted from a luxuriant prompt base in prior work [43], comprising prompt engineers’ prior knowledge; 2) Dynamic modifiers, which are extracted from the target showcase’s captions, and generated on the fly. They serve as a supplement to fill in more showcase-specific details dynamically.
2. Contextual Matching (Section IV-C). After obtaining the comprehensive modifier list, it is inefficient to query all of these modifiers to select the desired ones. Therefore, we refine them to narrow down the search space. Considering the modifiers’ decorative effect on the subject, we propose a contextual matching algorithm to refine these modifiers for streamlining the sample space. We verify that this intuitive matching method could achieve a more consistent ranking of the modifiers.
@@
3. Greedy Proxy Query (Section IV-D). Third, we optimize the prompts to achieve higher functionality. Since the local ranking from contextual matching may not fully align with the target, it is crucial to perform additional refinement based on feedback from the Proxy. We employ the greedy search algorithm to enhance the modifiers further. Specifically, for each prompt candidate, we query the Proxy and compute the similarity between its output and the showcase. Using the similarity gain as a criterion, we decide whether to retain or discard each modifier. This greedy approach allows us to identify the most effective modifiers. The entire interaction completes within $\mathcal{Q}$ turns.
@@ IV-B Modifier Extraction
To address this, we introduce two additional image elements, including subject detail and position information, to enrich the image description. In this paper, these are collectively referred to as dynamic modifiers, as they vary with each showcase.
Dynamic Modifier. We leverage the zero-shot capability of caption model to extract dynamic modifiers. Specifically, given a showcase, we first generate lots of prompts with a local BLIP model. This caption model could generate captions with large variances under a high sampling temperature, finding all the potential descriptions. Then, for each sampled caption, we utilize sentence analysis techniques to extract nominal and prepositional phrases. Specifically, we utilize Spacy [46] to analyze the prompt for Parts of Speech (PoS) and construct a parent-child tree for each word. For subject detail, we extract all the noun chunks in the prompt, including single nouns (e.g., “Black Panther”) and noun phrases (e.g., “red car”). For position information, the extraction process is much more complicated. First, Prometheus goes through one caption to locate all the preposition words. If one word is a preposition, then a prepositional phrase is constructed based on its relation tree. Take the phrase “during a sunset” as an example: when Prometheus detects that “during” is a preposition, the relation tree of “during” is extracted, e.g., “a” is the child of “sunset” and “a sunset” is the parent of “during”. Finally, we obtain a prepositional phrase by appending the child to the corresponding parent. We do not adopt an object detector as it can only recognize the category of a subject (e.g., car) while ignoring fine-grained details (e.g., the color of a car). Also, it cannot get the layout of an image, i.e., position information.
@@ IV-D Greedy Proxy Query
Score Function. To accurately assess the fidelity gain of a modifier, a comprehensive scoring function is essential. Intuitively, the perception of similarity between the recovered image (i.e., the image generated using the stolen prompt) and the showcase can serve as a useful guide. However, because the showcase includes many random elements introduced by the inherent randomness of any T2I model, relying too heavily on perception consistency risks severe overfitting. Specifically, the stolen prompt may end up including additional modifiers beyond the ground truth to account for this randomness. To address this, we leverage both semantic and perceptual guidance to refine the modifiers. Let $p_{b}$ represent the base prompt, which is updated to $p_{q}$ with modifier $m_{q}$, and let the corresponding feedback from the Proxy $\mathcal{O}$ be denoted as $\mathcal{O}(p_{b})$ and $\mathcal{O}(p_{q})$, respectively. The fidelity gain is:
@@
Greedy Proxy Query. We present our $\mathtt{GreedyProxyQuery}$ process in Algorithm 1, which contains four main steps:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
$\hat{p}=p_{b}+m_{d}+m_{s}.$ (3)
@@ V-A Experiment Setup
Datasets. We consider two datasets: DALLEPrompt and RealPrompt. Each data point in these sets consists of a showcase and the corresponding prompt.
@@
Baselines. We consider the only prompt stealing method and four prompt recovery methods as baselines. The content in brackets denotes the original task.
@@
    PH2P (Prompt Inversion[19]: We also consider PH2P, a method that inverts the prompt with white-box access to the generation model. PH2P first performs token-level optimization on the soft prompt using gradients, and then projects the optimized soft prompt into a hard prompt.
@@
Evaluation Metrics. We consider four metrics to evaluate the stolen prompt in terms of image semantic similarity, image perceptual similarity, prompt semantic similarity, and attack success rate:
@@
    LPIPS: The Learned Perceptual Image Patch Similarity (LPIPS) metric is designed to align closely with human perception [53]. It evaluates the visual similarity between two images using features extracted by a deep neural network (e.g., an AlexNet-based backbone). We use this metric to assess the perceptual similarity between the recovered image and the showcase.
@@
Besides these metrics, we also conducted a large-scale user study, which directly evaluates the human perception of the stolen effect. The detailed experimental design and results are presented in Section V-D.
Implementation Details. We implement Prometheus using Python 3.8 with PyTorch. All experiments are conducted on a single NVIDIA GeForce RTX A6000 GPU. The fidelity gain threshold in Equation 1 is set to 0.005. For BLIP used across all methods, we utilize the official implementation and adopt a ViT-based backbone, which is fine-tuned on the COCO dataset [17]. Unless otherwise specified, the Oracle query limit for Prometheus is set to 200, and the number of captions is set to 400. Note that, since we employ a local proxy model for feedback, the money cost is 0.
V-B Main Result
Functionality. Table II shows the main experimental results. In general, Prometheus surpasses all the baselines on both DALLEPrompt and RealPrompt, demonstrating its functionality and practicality in real-world scenarios. Specifically, it achieves the best average scores: image semantic similarity (0.897), image perception similarity (0.601), prompt semantic similarity (0.803), and ASR (56.0%) on DALLEPrompt. Similarly, on RealPrompt, Prometheus leads with image semantic similarity (0.912), image perception similarity (0.625), prompt semantic similarity (0.814), and ASR (62.5%), demonstrating its effectiveness in real-world scenarios.
For image semantics and perception, Prometheus’s improvements are mainly attributed to its prompt optimization with feedback from Proxy, while other methods only leverage local models to predict the prompt. This feedback-based refinement helps Prometheus generate images that are perceptually similar to the showcase. For prompt semantics, Prometheus also achieves better results on all datasets and proxy models. The main reason is that we improve the quality of recovered prompts by introducing dynamic modifiers and contextual matching. The former adds subject details as well as position information, and the latter ensures the sorted modifier achieves more consistent fidelity improvement when being put into the base prompt. Therefore, Prometheus obtains higher prompt semantic similarity with the victim prompt.
@@ V-B Main Result
Figure 6 illustrates eight examples of the attack effect achieved by Prometheus compared to baseline methods. Additionally, Table III presents an example of the stolen prompts. Overall, compared to the baselines, Prometheus generates prompts and images with higher fidelity to the ground truths and showcases in terms of image semantics, visual perception, and prompt semantics. Take Figure 6 case VI as an example. The baseline BLIP struggles to recognize the style (e.g., “hyperrealistic”) and fails to capture the main content of the image (e.g., “in a concert hall”). Meanwhile, CLIP-IG tends to match irrelevant modifiers to the showcase and introduces excessive redundancy. For instance, it ranks “3D cartoon” as having the highest correlation to the showcase (VI) and includes many unrelated “Disney” modifiers. These mismatches result in semantically incorrect image content and lower perceptual quality. This behavior can be attributed to CLIP-IG treating modifiers independently, leading to poorly contextualized predictions. PromptStealer exhibits two notable shortcomings. First, it demonstrates lower fidelity when predicting the subject (e.g., Mickey Mouse in case II and The Simpsons in case VII). Second, as shown in Table III, PromptStealer often predicts generic modifiers that are frequent in its training data [43], rather than contextually appropriate ones. A likely explanation is that PromptStealer’s prediction models are trained on out-of-domain data, such as that collected from Stable Diffusion, leading to overfitting. As for PH2P [19], this method fails to capture style modifiers and often even the subject. This highlights a fundamental limitation of using gradient-based prompt inversion for prompt stealing, indicating the need for further refinement to enable adaptation. Another baseline, VLMasExpert, performs relatively well but frequently omits key characters in the showcase. This limitation is likely due to alignment constraints during the model training and the out-of-distribution problem. We provide a more detailed analysis about the impact of alignment on its stealing performance in Appendix A-A. In contrast, Prometheus captures dynamic and contextually relevant modifiers, such as “auditorium” and “orchestra”, which align with elements present in the showcase. As we can see, in the base prompt, Prometheus also ignores the “concert hall” element, but the dynamic modifier fixes this issue. Furthermore, Prometheus accurately identifies the modifier “hyper realism”, demonstrating its outstanding prediction ability.
@@
Reusability on Different Subjects. Table IV presents the comparison of reusability between Prometheus and baseline methods. Reusability is a crucial property, as it enables users to modify the subject of a prompt while maintaining similar styles in the generated images. To evaluate this, we use the RealPrompt dataset, which comprises 8 style sets (three subjects per style) to simulate real-world scenarios. For each stolen prompt, an LLM is employed to replace the subject, producing a new prompt with a different subject. This process is achieved using the following instruction with GPT-4o: “Please replace the central subject in the sentence prompt with subject while keeping the rest of the content unchanged”. We validate the reusability of Prometheus and other baselines from both qualitative and quantitative angles. We present the experimental results in Table IV and visualization in Figure 7.
Multiple Showcases. We consider a practical scenario where multiple showcases are provided under a single theme. In this case, the attacker extracts a prompt for each showcase, resulting in a set of prompts. These prompts are then merged into a distilled prompt. Here, we consider using a large language model (i.e., ChatGPT-4o) to summarize the stolen prompts with the instruction “Please summarize these sentences into one sentence by rephrasing them. Replace $\left\{subjects\right\}$ or related words with $\left\{target\ subject\right\}$”. The experimental results are presented in Table V, leading to two key observations. First, while Prometheus continues to outperform the baseline methods, the margin of improvement is less pronounced. A potential explanation is that the baseline methods compensate for inadequacies in single stolen prompts by leveraging the collective information from the entire prompt set, thereby reducing Prometheus’s relative advantage. Second, compared to the results in Table II, the prompt semantic fidelity of Prometheus improves (from 0.814 to 0.840) after merging all the stolen prompts. This is likely because the summarization process mitigates overfitting to individual showcases, resulting in a stolen prompt that is better aligned with the ground truth. Take-away: When given multiple showcases on the same theme, Prometheus outperforms the baselines and achieves better performance than in single-showcase scenarios.

### v2

@@ I Introduction
Responsible Disclosure:
We obtained prior consent from both the prompt engineers and the platforms for the scientific use of the displayed prompt showcases and the collected data. They have expressed great interest in and support for our work. Upon acceptance, we will share the research findings with them, along with mitigation methods for such attacks. We will open-source the code and collected dataset for community use.
@@ II-B Prompt-as-a-Service
Despite the remarkable generation ability of T2I models, performing prompt engineering requires a significant amount of effort and cost. For instance, PromptPerfect charges up to \$100 per month for its prompt optimization service [51]. Under these circumstances, Prompt-as-a-Service (PaaS) has emerged as a popular application. Professional prompt engineers optimize prompts with their expertise and sell them on the prompt markets, e.g., PromptBase [48], AIFrog [4], PromptSea [53], PromptHero [50]. These on-sell prompts are sightlessly exhibited with their exquisite showcases, which visually indicate the prompt effects. After purchasing, the user gains access to its corresponding textual description, which can be used to generate customized images or increase the number of showcased outputs.
@@ III-A Threat Model
Attacker’s Capability.
We make two assumptions about the attacker’s capabilities. (1) The attacker can access the generated showcases of the prompt (denoted as $s$). This is typical in commercial platforms, where prompt sellers display showcase images to attract potential buyers. However, the specific prompt used to generate the showcase remains hidden from all users until they purchase it. A detailed example can be found in Portrait Neon Lights [52]. (2) As some commercial victim models run in a pay-as-you-go mode, which may be expensive, we assume the attacker can use a substitute local text-to-image (T2I) model, referred to as the proxy model. Since the proxy model can be any open-source model, the query budget is 0.
Adversary’s Motivation.
The attacker aims to compromise the confidentiality of valuable prompts sold on the markets. A carefully designed prompt consists of subjects and modifiers (as shown in Figure 1). Specifically, subjects are the main body of a prompt that outlines an image’s most important elements, while modifiers detail the style of the subjects and the whole picture. Formally, given a showcase image $s$, an adversary aims to recover the prompt $p$ that is used to create the showcase, including both its subjects and modifiers. We provide more analysis about the motivation in Appendix A-A.
From a utility perspective, stolen prompts can be reused for customization, e.g., replacing subjects or increasing the number of showcases. From a commercial standpoint, as discussed in previous work [68], the motivation behind prompt stealing can be twofold. First, it may infringe on the intellectual property (IP) of high-quality prompts, which is the result of significant effort by prompt engineers. Second, attackers have strong economic incentives: they can avoid payment by using these premium prompts for free and may also redistribute or resell the prompts on other markets for additional profit. This poses a significant threat to the commercial interests of both prompt developers and the platforms hosting them, which is explicitly prohibited by platform PromptBase [49]. From a privacy protection standpoint, this type of attack poses a significant threat to the confidentiality of prompts, particularly when they contain sensitive or highly classified information. We provide more analysis for this standpoint in Appendix A-B.
@@ III-B Attack Requirements
Attack Comparisons.
We follow the existing line of work [27, 47, 68], and focus on reversing readable prompts, which could satisfy these two objectives. We observe that certain hard prompt recovery methods could serve as potential solutions. For instance, PH2P [37] leverages gradient signals to optimize a soft prompt, which is then projected back into a hard prompt. However, since semantic coherence is not taken into account during optimization, the recovered prompts often fall into a local optimum and consist of unreadable or nonsensical words. Another drawback of hard prompt recovery methods is that they search modifier candidates within the entire discrete token space, which is rather inefficient. For example, when recovering a 40-word prompt from a 50,000-token vocabulary, PH2P would consider $50,000^{40}$ combinations, which is computationally infeasible. As a result, they fail to meet the aforementioned objectives, limiting the practical utility of the extracted prompts. We list the merits of our Prometheus over the existing works in Table I. Besides static modifiers, we also include dynamic modifiers and feedback from the proxy model, to steal prompts with higher fidelity from a meaningful modifier set efficiently.
@@ IV-A Overview
1. Modifier Extraction
(Section IV-B). First, we construct a comprehensive and semantic modifier pool to ensure functionality and reusability. We satisfy these requirements by considering two kinds of modifiers: 1) Static modifiers, which are extracted from a luxuriant prompt base in prior work [68], comprising prompt engineers’ prior knowledge; 2) Dynamic modifiers, which are extracted from the target showcase’s captions, and generated on the fly. They serve as a supplement to fill in more showcase-specific details dynamically.
2. Contextual Matching
(Section IV-C). After obtaining the comprehensive modifier list, it is inefficient to query all of these modifiers to select the desired ones. Therefore, we refine them to narrow down the search space. Considering the modifiers’ decorative effect on the subject, we propose a contextual matching algorithm to refine these modifiers for streamlining the sample space. We verify that this intuitive matching method could achieve a more consistent ranking of the modifiers.
@@ 2. Contextual Matching
3. Greedy Proxy Query
(Section IV-D). Third, we optimize the prompts to achieve higher functionality. Since the local ranking from contextual matching may not fully align with the target, it is crucial to perform additional refinement based on feedback from the Proxy. We employ the greedy search algorithm to enhance the modifiers further. Specifically, for each prompt candidate, we query the Proxy and compute the similarity between its output and the showcase. Using the similarity gain as a criterion, we decide whether to retain or discard each modifier. This greedy approach allows us to identify the most effective modifiers. The entire interaction completes within $\mathcal{Q}$ turns.
@@ IV-B Modifier Extraction
To address this, we introduce two additional image elements, including subject detail and position information, to enrich the image description. In this paper, these are collectively referred to as dynamic modifiers, as they vary with each showcase. Formally, the dynamic modifiers are defined as the subject details (e.g., identity and specified property) and the position information (i.e., the whole layout of showcase), which are beyond the static style modifiers.
Dynamic Modifier.
We present the pseudocode (Appendix A-C) along with the following introduction to detail dynamic modifier extraction. We leverage the zero-shot capability of caption model to extract dynamic modifiers. Specifically, given a showcase, we first generate multiple prompts with a local BLIP model. This caption model could generate captions with large variances under a high sampling temperature, finding all the potential descriptions. Then, for each sampled caption, we utilize sentence analysis techniques to extract nominal and prepositional phrases. Specifically, we utilize Spacy [71] to analyze the prompt for Parts of Speech (PoS) and construct a parent-child tree for each word. For subject detail, we extract all the noun chunks in the prompt, including single nouns (e.g., “Black Panther”) and noun phrases (e.g., “red car”). For position information, the extraction process is much more complicated. First, Prometheus goes through one caption to locate all the preposition words. If one word is a preposition, then a prepositional phrase is constructed based on its relation tree. Take the phrase “during a sunset” as an example: when Prometheus detects that “during” is a preposition, the relation tree of “during” is extracted, e.g., “a” is the child of “sunset” and “a sunset” is the parent of “during”. Finally, we obtain a prepositional phrase by appending the child to its corresponding parent. We do not adopt an object detector as it can only recognize the category of a subject (e.g., car) while ignoring fine-grained details (e.g., the color of a car). Also, it cannot get the layout of an image, i.e., position information.
@@ IV-D Greedy Proxy Query
Score Function.
To accurately assess the fidelity gain of a modifier, a comprehensive scoring function is essential. Intuitively, the perception of similarity between the recovered image (i.e., the image generated using the stolen prompt) and the showcase can serve as a useful guide. However, because the showcase includes many random elements introduced by the inherent randomness of any T2I model, relying too heavily on perception consistency risks severe overfitting. Specifically, the stolen prompt may end up including additional modifiers beyond the ground truth to account for this randomness. To address this, we leverage both semantic and perceptual guidance to refine the modifiers. Let $p_{b}$ represent the base prompt, which is updated to $p_{q}$ with modifier $m_{q}$, and let the corresponding feedback from the Proxy $\mathcal{O}$ be denoted as $\mathcal{O}(p_{b})$ and $\mathcal{O}(p_{q})$, respectively. The fidelity gain is:
@@
Greedy Proxy Query.
We present our $\mathtt{GreedyProxyQuery}$ process in Algorithm 1, which contains four main steps:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
$\hat{p}=p_{b}+m_{d}+m_{s}.$ (3)
@@ V-A Experiment Setup
Datasets.
We consider two datasets: DALLEPrompt and RealPrompt. Each data point in these sets consists of a showcase and the corresponding prompt.
@@
Baselines.
We consider the only prompt stealing method and four prompt recovery methods as baselines. The content in brackets denotes the original task.
@@
    PH2P (Prompt Inversion[37]: We also consider PH2P, a method that inverts the prompt with access to the encoder of Stable Diffusion v1.5. PH2P first performs token-level optimization on the soft prompt using gradients, and then projects the optimized soft prompt into a hard prompt.
@@
Evaluation Metrics.
We consider four metrics to evaluate the stolen prompt, including:
@@
    LPIPS: The Learned Perceptual Image Patch Similarity (LPIPS) metric is designed to align closely with human perception [84]. It evaluates the visual similarity between two images using features extracted by neural network (e.g., AlexNet backbone). We use LPIPS to assess the perceptual similarity between the recovered image and showcase. A lower LPIPS indicates better perception alignment.
@@
We agree that no single metric fully captures similarity or usefulness, and selecting the top image by one metric can conceal weaknesses on others. Accordingly, for the three attacks, we do not pick a single “best” test by any one metric; instead, we report the average across all metrics. Besides these metrics, we also conducted a large-scale user study, which directly evaluates the human perception of the stolen effect. The detailed experimental design and results are presented in Section V-D.
@@
Implementation Details.
We implement Prometheus using Python 3.8 with PyTorch. All experiments are conducted on a single NVIDIA GeForce RTX A6000 GPU. The fidelity gain threshold in Equation 1 is set to 0.005. For BLIP used across all methods, we utilize the official implementation and adopt a ViT-based backbone, which is fine-tuned on the COCO dataset [33]. Unless otherwise specified, the Oracle query limit for Prometheus is set to 200, and the number of captions is set to 400. Note that, since we employ a local proxy model for feedback, the money cost is 0. To tackle the randomness within the attack process and image generation, we run each experiment three times.
V-B Main Result
Functionality.
Table II shows the main experimental results. In general, Prometheus surpasses all the baselines on both DALLEPrompt and RealPrompt, demonstrating its functionality and practicality in real-world scenarios. Specifically, it achieves the best average scores: image semantic similarity (0.897), image perception similarity (0.601), prompt semantic similarity (0.803), and ASR (56.0%) on DALLEPrompt. Similarly, on RealPrompt, Prometheus leads with image semantic similarity (0.912), image perception similarity (0.625), prompt semantic similarity (0.814), and ASR (62.5%), demonstrating its effectiveness in real-world scenarios.
@@ V-B Main Result
For image semantics and perception, Prometheus’s improvements are mainly attributed to its prompt optimization with feedback from Proxy, while other methods only leverage local models to predict the prompt. This feedback-based refinement helps Prometheus generate images that are perceptually similar to the showcase. For prompt semantics, Prometheus also achieves better results on all datasets and proxy models. The main reason is that we improve the quality of recovered prompts by introducing dynamic modifiers and contextual matching. The former adds subject details as well as position information, and the latter ensures the sorted modifier achieves more consistent fidelity improvement when being put into the base prompt. Therefore, Prometheus obtains higher prompt semantic similarity with the victim prompt. We include a more detailed analysis of the performance gap between DALLEPrompt and RealPrompt in Appendix B-D.
@@
Figure 6 illustrates eight examples of the attack effect achieved by Prometheus compared to baseline methods. Additionally, Table III presents an example of the stolen prompts. Overall, compared to the baselines, Prometheus generates prompts and images with higher fidelity to the ground truths and showcases in terms of image semantics, visual perception, and prompt semantics. Take Figure 6 case VI as an example. The baseline BLIP struggles to recognize the style (e.g., “hyperrealistic”) and fails to capture the main content of the image (e.g., “in a concert hall”). Meanwhile, CLIP-IG tends to match irrelevant modifiers to the showcase and introduces excessive redundancy. For instance, it ranks “3D cartoon” as having the highest correlation to the showcase (VI) and includes many unrelated “Disney” modifiers. These mismatches result in semantically incorrect image content and lower perceptual quality. This behavior can be attributed to CLIP-IG treating modifiers independently, leading to poorly contextualized predictions. PromptStealer exhibits two notable shortcomings. First, it demonstrates lower fidelity when predicting the subject (e.g., Mickey Mouse in case II and The Simpsons in case VII). Second, as shown in Table III, PromptStealer often predicts generic modifiers that are frequent in its training data [68], rather than contextually appropriate ones. A likely explanation is that PromptStealer’s prediction models are trained on out-of-domain data, such as that collected from Stable Diffusion, leading to overfitting. As for PH2P [37], this method fails to capture style modifiers and often even the subject. This highlights a fundamental limitation of using gradient-based prompt inversion for prompt stealing, indicating the need for further refinement to enable adaptation. Another baseline, VLMasExpert, performs relatively well but frequently omits key characters in the showcase. This limitation is likely due to alignment constraints during the model training and the out-of-distribution problem. We provide a more detailed analysis about the impact of alignment on its stealing performance in Appendix B-A. In contrast, Prometheus captures dynamic and contextually relevant modifiers, such as “auditorium” and “orchestra”, which align with elements present in the showcase. As we can see, in the base prompt, Prometheus also ignores the “concert hall” element, but the dynamic modifier fixes this issue. Furthermore, Prometheus accurately identifies the modifier “hyper realism”, demonstrating its outstanding prediction ability.
@@
Reusability on Different Subjects.
Table IV presents the comparison of reusability between Prometheus and baseline methods. Reusability is a crucial property, as it enables users to modify the subject of a prompt while maintaining similar styles in the generated images. To evaluate this, we use the RealPrompt dataset, which comprises 8 style sets (three subjects per style) to simulate real-world scenarios. For each stolen prompt, an LLM is employed to replace the subject, producing a new prompt with a different subject. This process is achieved using the following instruction with GPT-4o: “Please replace the central subject in the sentence prompt with subject while keeping the rest of the content unchanged”. We validate the reusability of Prometheus and other baselines from both qualitative and quantitative angles. We present the experimental results in Table IV and visualization in Figure 7.
@@
Multiple Showcases.
We consider a practical scenario where multiple showcases are provided under a single theme. In this case, the attacker extracts a prompt for each showcase, resulting in a set of prompts. These prompts are then merged into a distilled prompt. Here, we consider using a large language model (i.e., ChatGPT-4o) to summarize the stolen prompts with the instruction “Please summarize these sentences into one sentence by rephrasing them. Replace $\left\{subjects\right\}$ or related words with $\left\{target\ subject\right\}$”. The experimental results are presented in Table V, leading to two key observations. First, while Prometheus continues to outperform the baseline methods, the margin of improvement is less pronounced. A potential explanation is that the baseline methods compensate for inadequacies in single stolen prompts by leveraging the collective information from the entire prompt set, thereby reducing Prometheus’s relative advantage. Second, compared to the results in Table II, the prompt semantic fidelity of Prometheus improves (from 0.814 to 0.840) after merging all the stolen prompts. This is likely because the summarization process mitigates overfitting to individual showcases, resulting in a stolen prompt that is better aligned with the ground truth. Take-away: When given multiple showcases on the same theme, Prometheus outperforms the baselines and achieves better performance than in single-showcase scenarios.